a sales person earns a basic salary of $600 per month plus commission of 10% on sales she makes. she finds that on average she takes 1 1/2 hours to make $100 worth of sales. how many hours must she work on the average each month if her monthly earnings are to be $2000?
Answer by ejazali_syed(7)

Q) a sales person earns a basic salary of $600 per month plus commission of 10% on sales she makes. she finds that on average she takes 1 1/2 hours to make $100 worth of sales. how many hours must she work on the average each month if her monthly earnings are to be $2000?
A) the sales person earns basic salary plus commission on the salary that sums up to his total salary.

if the basic salary is x; his/her total salary would be
x + x*10% = T where T is the total salary for a month.
given T = 2000
x + x*10% = 2000 => x + 0.1x = 2000
==> 1.1X = 2000 ==> x = 2000/1.1 = 1818.18
hence the sales that person has made is 1818.18

if it takes 1.5 hrs to make $100 worth of sales.
then ??? (T)hrs to make 1818.18 worth of sales??
hence T = 1818.18*1.5/100 = 27.2727 hrs